Solve the following problem. It may be helpful to use draw a chart on scrap paper to organize the information and write the equation. Be sure to show all steps (V.E.S.T.) and work in order to receive full credit.


A grocer wants to make a 10-pound mixture of cashews and peanuts that he can sell for $3.29 per pound. If cashews cost $5.60 per pound and peanuts cost $2.30 per pound, how many pounds of each must he mix?

Answer by stanbon(75887)

A grocer wants to make a 10-pound mixture of cashews and peanuts that he can sell for $3.29 per pound. If cashews cost $5.60 per pound and peanuts cost $2.30 per pound, how many pounds of each must he mix?
---------------------
Using 2 equations:
Quantity: c + p = 10 lbs
Value::: 5.6c + 2.3p = 3.29*10 dollars
----------
Multiply thru the Quantity Eq. by 560
Multiply thru the Value Eq. by 100
-------------------
560c + 560p = 5600
560c + 230p = 3290
-----
Subtract and solve for "p":
330p = 2310
p = 7 lbs (amt. of peanuts needed)
------
Solve for "c" using c+p = 10
c + 7 = 10
c = 3 lbs (amt. of cashews needed)
=====================================
cheers,
Stan H.